Rensing Ltd. estimates sales for the second quarter of 2017 will be as follows.

Month

Units

April 2,520
May 2,450
June 2,350


The target ending inventory of finished products is as follows.

March 31 2,030
April 30 2,220
May 31 2,110
June 30 2,360


2 units of material are required for each unit of finished product. Production for July is estimated at 2,640 units to start building inventory for the fall sales period. Rensing’s policy is to have an inventory of raw materials at the end of each month equal to 60% of the following month’s production requirements.

Raw materials are expected to cost $4 per unit throughout the period.

Calculate the May raw materials purchases in dollars.

Raw material purchases cost

$   
Answer #1

May production Unit = 2450+2110-2220 = 2340 Units

June Production Unit = 2350+2360-2110 = 2600 Units

Raw material purchase = (2340*2)+(2600*2*60%)-(2340*2*60%) = 4992

Raw material purchase cost = 4992*4 = $19968